OpenStudy (mantar0078):

Could someone give me the steps to answer this? I can't remember them. Find A + B, if possible. A=[1.5 3.8 3] [-1.2 2.4 0] B=[0 4 1] [0 -2 1]

OpenStudy (triciaal):

please verify the post

OpenStudy (mantar0078):

What do you mean?

OpenStudy (triciaal):

verify means to check and confirm that what you are verifying is correct

OpenStudy (triciaal):

what you have posted is not clear as posted

OpenStudy (sshayer):

\[A+B=\left[\begin{matrix}1 & 53&83 \\ -1 & 22 &40\end{matrix}\right]+\left[\begin{matrix}0 & 4&1 \\ 0 & -2&1\end{matrix}\right]\] \[=\left[\begin{matrix}1+0 & 53+4&83+1 \\ -1+0 & 22-2&40+1\end{matrix}\right]=?\]

OpenStudy (sshayer):

i assumed you are asking about the sum of two matrices A and B

OpenStudy (mantar0078):

Yes I was. Thank you sshayer

OpenStudy (triciaal):

@sshayer good you made that assumption was no where near my mind was confused about regular sets

OpenStudy (sshayer):

you both are welcome.